# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v3

@@ Results
After the removal of $285$ individuals due to missing covariate information, the analysis cohort consisted of $3489$ HCT study participants with positive SARS-CoV-2 tests prior to February $24$, $2023$. Of these, $1434$ ($41.1$%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was $23$ years (IQR: 2020202036363636), and $64$% of the participants reported female sex. Slightly over $33$% of participants reported at least one baseline comorbidity, and roughly $52$% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at $30$, $60$, and $90$ days after the positive test were $19.2$% (95% CI: 13.713.713.713.725.225.225.225.2), $12.9$% (9.49.49.49.416.516.516.516.5), and $6.8$% (2.82.82.82.810.910.910.910.9), respectively.
@@
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than $1$ indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of $1.21$ (1.041.041.041.041.421.421.421.42, $P=0.014$). Conversely, participants reporting fatigue at baseline tended to have longer lasting symptoms, with an estimated hazard ratio of $0.77$ (0.650.650.650.650.930.930.930.93, $P=0.005$). In terms of viral load, participants with Ct values under $30$ and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over $30$, with estimated hazard ratios of $0.78$ (0.640.640.640.640.960.960.960.96) and $0.76$ (0.600.600.600.600.960.960.960.96), respectively ($P=0.034$).
@@ Discussion
We demonstrated the use of CIR, along with existing survival analysis regression methods, to analyze current status data from the long COVID survey study conducted at UW between December $2021$ and June $2023$. Participants were queried at least $28$ days after testing positive, and even for those queried exactly at $28$ days, many did not respond until days or weeks later. This variation in response times motivates the use of the statistical methods described in this paper. Using the long COVID survey data, we estimated the prevalence of persistent COVID symptoms in the UW population to be $19$% at $30$ days after a positive test, dropping to $7$% at $90$ days.
A retrospective study in a comparable university population found that $36$% of survey respondents reported symptoms lasting at least 28 days from the end of the recommended $10$-day isolation period (Landry et al.,, 2023). Our results provide further evidence that, even in a predominantly young and vaccinated population, a substantial proportion of infected individuals are affected by persistent COVID-19 symptoms. Notably, because participants were asked only about persistent symptoms and not about new symptoms arising in the post-acute phase, our analysis may still represent an underestimate of the frequency of PASC, which may also be characterized by symptoms that differ from initial acute symptoms and arise after their resolution.
We also investigated potential risk factors for persistent symptoms using Cox proportional hazards regression, and found evidence that female sex, fatigue during the acute phase, and higher semi-quantitative viral load at time of testing are associated with slower resolution of symptoms. This is in line with previous studies that have identified female sex (Su et al.,, 2022), severe fatigue during the acute infection phase (Shen et al.,, 2023), and higher peak semi-quantitative viral load in the four weeks following diagnosis (Lu et al.,, 2023) as potential risk factors for PASC.
@@ Acknowledgements
Without making strong modeling assumptions, the target regression function $\theta_{0}(y_{0})$ is not pathwise differentiable and is therefore not estimable at $n^{-\frac{1}{2}}$–rate. However, it is known to be monotone as a function of $y_{0}$ and can be viewed as the target of an isotonic regression procedure. Using isotonic regression to estimate $\theta_{0}(y_{0})$ allows for the application of established statistical theory (Westling and Carone,, 2020) to characterize the large-sample properties of the resulting estimator. We outline this isotonic regression procedure here.
For a generic $P\in\mathcal{M}$, we first define the outcome regression function $\mu_{P}:(y,w)\mapsto E_{P}(\Delta\,|\,W=w,Y=y)$ and density ratio $g_{P}:(y,w)\mapsto\pi_{P}(y\,|\,w)/f_{P}(y)$ with $\pi_{P}$ the conditional density of $Y$ given $W$ and $f_{P}$ the marginal density of $Y$. These nuisance functions play a key role in accounting for the possible dependence between $\Delta_{*}$ and $Y_{*}$. We use $F_{P}$ to denote the marginal distribution function of $Y$. We use $F_{P}$ as a domain transformation function in the CIR procedure. While the CIR procedure could be performed without the use of this domain transformation, its use results in a procedure that is invariant to strictly increasing transformations of $Y$ (Westling et al.,, 2020). In other words, changing the scale on which response times are measured will have no impact on the estimated survival function.
[…]
@@
with $Q_{P}$ the marginal distribution function of $W$ under sampling from $P$. Unlike $\theta_{0}(y_{0})$, the parameter $\Gamma_{0}(y_{0})$ is pathwise differentiable relative to the nonparametric model $\mathcal{M}$ and can therefore be estimated at $n^{-\frac{1}{2}}$–rate (Pfanzagl and Pfanzagl,, 1990). Per the results of Westling et al., (2020), the efficient influence function of $P\mapsto\Gamma_{P}(y_{0})$ relative to $\mathcal{M}$ evaluated at $P_{0}$ is given by
@@
Knowledge of this efficient influence functions enables standard debiasing approaches, such as the one-step correction (Pfanzagl,, 1982), to be used in estimation of $\Gamma_{0}(y_{0})$. This yields a debiased, asymptotically linear estimator $\Gamma_{n}(y_{0})$ of $\Gamma_{0}(y_{0})$, which is used in the isotonic regression procedure to construct an estimator of $\theta_{0}(y_{0})$. Our proposed procedure is as follows.
@@
    Select domain bounds $t_{0}$ and $t_{1}$ satisfying $[t_{0},t_{1}]\subset(b_{0},c_{0})$.
@@
We note that, were we not to apply the domain transformation $F_{P}$, we could perform a similar procedure using the primitive parameter $y_{0}\mapsto\iint\mathbbm{1}(u\leq y_{0})\mu_{P}(u,w)Q_{P}(dw)$. In the absence of survey nonresponse, this procedure would take the same form as the isotonization method of van der Vaart and van der Laan, (2006). However, it is not clear whether this approach can be framed in terms of isotonic regression using pseudo-outcomes, as described in the main text.
If for every point $y_{0}$ in the estimation domain $[t_{0},t_{1}]$, either $\mu_{n}$ converges to $\mu_{0}$ or $g_{n}$ converges to $g_{0}$, Theorem 1 of Westling et al., (2020) establishes that $\theta_{n}(y_{0})$ converges in probability to $\theta_{0}(y_{0})$ for each $y_{0}\in[t_{0},t_{1}]$. The conditions used to establish distributional convergence are more complex. Roughly speaking, we require that the product of the convergence rates of $\mu_{n}$ and $g_{n}$ to their respective targets is $n^{-\frac{1}{3}}$. If these conditions are satisfied, Theorem 2 of Westling et al., (2020) establishes that, for any $y_{0}\in(b_{0},c_{0})$ with $F_{0}(y_{0})\in(0,1)$, $n^{\frac{1}{3}}\left\{\theta_{n}(y_{0})-\theta_{0}(y_{0})\right\}$ converges in distribution to $\tau_{0}(y_{0})^{\frac{1}{3}}\mathbb{W}$, where $\mathbb{W}$ follows a standard Chernoff distribution (Groeneboom and Wellner,, 2001) and $\tau_{0}(y_{0}):=4\theta_{0}^{\prime}(y_{0})\kappa_{0}(y_{0})/f_{0}(y_{0})$ with
@@
A copula is a multivariate cumulative distribution function, indexed by an association parameter $\alpha$, that describes the joint distribution of two or more random variables, each with a uniform marginal distribution. Sklar’s theorem (Sklar,, 1959) states that any multivariate distribution can be expressed in terms of its marginal distributions and a copula. Throughout this section, we use $F$ and $f$ to denote generic distribution and density functions, respectively, with a subscript indicating the distribution and random variable corresponding to the given distribution/density function, e.g., $F_{P;T\,|\,W}(t\,|\,w):=P(T\leq t\,|\,W=w)$.
@@
[…]
@@
$(\delta,y,w)\mapsto$ $\frac{\left(1-e^{-\alpha}\right)\mathbbm{1}(y\leq y_{0})\left\{\delta-\mu_{0}(y,w)\right\}e^{-\alpha H_{0}(y,w)}}{\alpha g_{0}(y,w)r_{\alpha}\{\mu_{0}(y,w),H_{0}(y,w)\}}-2\int_{u\leq y_{0}}\lambda_{0;\alpha}(u)F_{0}(du)$
@@
Many popular regression models used in survival analysis have been adapted for use with current status data, including the Cox proportional hazards model (Huang,, 1996), the proportional odds model (Rossini and Tsiatis,, 1996), and the accelerated failure time model (Rabinowitz et al.,, 1995). These methods are generally valid under the same assumptions on the survey response process as the adaptation of the CIR method described above. They do, however, make additional assumptions on the relationship between $T$ and $W$.
Fitting regression models using current status data is generally more complicated than with right censored data. For the Cox model, unlike in the right-censored setting, estimation of the regression coefficients cannot be completely separated from estimation of the baseline hazard function through the partial likelihood (Huang,, 1996). Nonetheless, estimation and valid inference is still attainable for the regression coefficients, with the main difficulties being computational in nature. Recent work has focused on developing more efficient algorithms for computing estimates of the regression coefficients via maximum likelihood (Anderson-Bergman,, 2017). Calculation of the observed Fisher information, which is used to compute analytic standard error estimates, is generally impractical. The bootstrap has been proposed as an alternative approach (Pan,, 1999), although as of yet there is no theoretical proof of its validity in this context. In Section S4, we present results of a simulation study investigating the performance of the nonparametric bootstrap for construction of confidence intervals using current status data for inference on regression coefficients of a Cox model. Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance, while percentile bootstrap confidence intervals tend to be slightly anticonservative. Based on these findings, we make use of the Wald-type interval in the regression analyses we report.
As with the Cox model, the other regression models described above do not need to be altered to account for survey nonresponse, under the assumption that the regression model is correctly specified and includes all covariates that inform the response mechanism (Little and Rubin,, 2019).
@@
For each simulation replicate, the aim was to estimate the distribution function of $T$ over the interval $[t_{0},t_{1}]$ with $t_{0}=0.02$ (corresponding to the 5th percentile of observed response times) and $t_{1}=1.5$. The maximum follow-up time was set to $c_{0}\in\{2.1,1.8,1.65\}$. We refer to the settings corresponding to these values of $c_{0}$ as Scenarios 1, 2 and 3, respectively. Under these three choices of $c_{0}$, $t_{1}$ corresponded to the 90th, 95th, or 97.5th quantile of observed survey response times, respectively. This allowed us to explore whether performance of the complete case CIR or extended CIR procedures would degrade if $c_{0}$ was chosen such that nonresponse was more prevalent and $t_{1}$ was nearer the edge of the support of observed response times. For each value of $c_{0}$, the observed data consisted of $n$ independent replicates of $(W,Y,\Delta)$ with $Y:=\min(Y_{*},c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\mathbbm{1}(T\leq Y)$. We generated datasets of size $n\in\{500,1000,1500,2000\}$. For each simulation setting, we performed 500 simulation replicates.
The survival function was estimated (1) using only complete cases (i.e., survey respondents, or those with $Y_{*}<c_{0}$) with the existing CIR procedure described in (Westling et al.,, 2020), (2) using all participants with the extended CIR procedure proposed here, and (3) using the NPMLE, as implemented in the survival software package (Therneau,, 2022), computed only using complete cases. For the two CIR procedures, nominal 95% confidence intervals were constructed by estimating the scale factor and using the quantiles of the Chernoff distribution. The nuisance function $\mu$ was estimated using Super Learner (van der Laan et al.,, 2007) with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman,, 1991), generalized additive models (GAMs) (Hastie,, 2017), and random forests (Wright and Ziegler,, 2015). The conditional density $\pi$ was estimated using haldensify (Hejazi et al.,, 2022) with 10-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. For each simulation replicate, the survival function of $T$ was estimated at 100 time-points, evenly spaced on the quantile scale, between $t_{0}=0.02$ and $t_{1}=1.5$. For each setting and method, we conducted 1000 simulation replicates. Pointwise and integrated (over all 100 time-points) performance was quantified using the empirical bias and empirical confidence interval coverage.
@@
We used the extended CIR procedure, with several different combinations of nuisance estimators, to estimate the survival function of $T$. The conditional density $\pi$ was estimated using either (1) a misspecified parametric log-normal accelerated failure time model including only linear terms, or (2) haldensify, an implementation of a nonparametric approach based on the highly adaptive lasso (Hejazi et al.,, 2022). The outcome regression nuisance function $\mu$ was estimated using either (1) logistic regression, (2) GAMs (Hastie,, 2017), (3) gradient boosted trees (Chen and Guestrin,, 2016), or (4) Super Learner (van der Laan et al.,, 2007) with a library as described in the previous experiment. For comparison, we also included the true form for $g$ and $\mu$ (i.e., in lieu of estimating the nuisance parameters, we plugged in their known values). As long as either of the two nuisance estimators is sufficiently flexible to consistently estimate its target, we expect the extended CIR procedure to yield a consistent estimator. Furthermore, if the nuisance estimation rate is fast enough, as described in Section S1.2, we expect the resulting extended CIR estimator to not only be consistent, but to have a large-sample distribution that does not depend on the particular nuisance estimator employed. This large-sample distribution be identical as when the nuisance parameters are set to their true values. For each sample size, we conducted 500 simulations replicates; all 15 possible combinations of nuisance estimators were applied to each simulated dataset. We examined pointwise empirical bias and pointwise empirical variance. Furthermore, we examined the absolute difference between survival function estimates computed on each dataset, comparing (1) using known nuisance parameters versus using the haldensify–Super Learner combination, and (2) using the haldensify–Super Learner combination versus using the haldensify–gradient boosted trees combination. These two comparisons quantify the sensitivity of the proposed procedure to the choice of nuisance estimator when a flexible method is used in place of the ‘ideal’ method using known nuisance parameter values and when two distinct flexible methods are used, respectively.
@@
We generated datasets of size $n\in\{500,1000,2000,4000\}$. The maximum follow-up time $c_{0}$ was set to 1.65. We aimed to estimate the cumulative distribution function of $T$ over $[t_{0},t_{1}]$, with $t_{0}=0.02$ and $t_{1}=1.5$. The estimation procedure was as described in Section S4.1. We compared the performance of the extended CIR procedure using the copula-based one-step estimator $\Gamma_{P_{n},\alpha}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},\alpha,y_{0}}(Z_{i})$ to the performance of the extended CIR procedure using $\Gamma_{P_{n}}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},y_{0}}(Z_{i})$, which we expect to provide consistent estimation only under conditional independence. We refer to the latter as the ‘baseline’ extended CIR. We quantified performance in terms of empirical bias. For each simulation setting, we performed 500 simulation replicates.
@@
For data analyses and simulations Super Learner (van der Laan et al.,, 2007) was implemented with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman,, 1991), generalized additive models (Hastie,, 2017), random forests (Wright and Ziegler,, 2015), and gradient boosted trees (Chen and Guestrin,, 2016).

### v4

@@ Results
After the removal of $285$ individuals due to missing covariate information, the analysis cohort consisted of $3489$ HCT study participants with positive SARS-CoV-2 tests prior to February $24$, $2023$. Of these, $1434$ (41.141.1%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was $23$ years (IQR: 20203636), and 6464% of the participants reported female sex. Slightly over 3333% of participants reported at least one baseline comorbidity, and roughly 5252% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at $30$, $60$, and $90$ days after the positive test were 19.219.2% (95% CI: 13.713.725.225.2), 12.912.9% (9.49.416.516.5), and 6.86.8% (2.82.810.910.9), respectively.
@@
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than $1$ indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of $1.21$ (1.041.041.421.42, $P=0.014$). Conversely, participants reporting fatigue at baseline tended to have longer lasting symptoms, with an estimated hazard ratio of $0.77$ (0.650.650.930.93, $P=0.005$). In terms of viral load, participants with Ct values under $30$ and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over $30$, with estimated hazard ratios of $0.78$ (0.640.640.960.96) and $0.76$ (0.600.600.960.96), respectively ($P=0.034$).
@@ Discussion
We demonstrated the use of CIR, along with existing survival analysis regression methods, to analyze current status data from the long COVID survey study conducted at UW between December $2021$ and June $2023$. Participants were queried at least $28$ days after testing positive, and even for those queried exactly at $28$ days, many did not respond until days or weeks later. This variation in response times motivates the use of the statistical methods described in this paper. Using the long COVID survey data, we estimated the prevalence of persistent COVID symptoms in the UW population to be 1919% at $30$ days after a positive test, dropping to 77% at $90$ days.
A retrospective study in a comparable university population found that 3636% of survey respondents reported symptoms lasting at least 28 days from the end of the recommended $10$-day isolation period (Landry et al., 2023). Our results provide further evidence that, even in a predominantly young and vaccinated population, a substantial proportion of infected individuals are affected by persistent COVID-19 symptoms. Notably, because participants were asked only about persistent symptoms and not about new symptoms arising in the post-acute phase, our analysis may still represent an underestimate of the frequency of PASC, which may also be characterized by symptoms that differ from initial acute symptoms and arise after their resolution.
We also investigated potential risk factors for persistent symptoms using Cox proportional hazards regression, and found evidence that female sex, fatigue during the acute phase, and higher semi-quantitative viral load at time of testing are associated with slower resolution of symptoms. This is in line with previous studies that have identified female sex (Su et al., 2022), severe fatigue during the acute infection phase (Shen et al., 2023), and higher peak semi-quantitative viral load in the four weeks following diagnosis (Lu et al., 2023) as potential risk factors for PASC.
@@ Acknowledgements
Without making strong modeling assumptions, the target regression function $\theta_{0}(y_{0})$ is not pathwise differentiable and is therefore not estimable at $n^{-\frac{1}{2}}$–rate. However, it is known to be monotone as a function of $y_{0}$ and can be viewed as the target of an isotonic regression procedure. Using isotonic regression to estimate $\theta_{0}(y_{0})$ allows for the application of established statistical theory (Westling and Carone, 2020) to characterize the large-sample properties of the resulting estimator. We outline this isotonic regression procedure here.
For a generic $P\in\mathcal{M}$, we first define the outcome regression function $\mu_{P}:(y,w)\mapsto E_{P}(\Delta\,|\,W=w,Y=y)$ and density ratio $g_{P}:(y,w)\mapsto\pi_{P}(y\,|\,w)/f_{P}(y)$ with $\pi_{P}$ the conditional density of $Y$ given $W$ and $f_{P}$ the marginal density of $Y$. These nuisance functions play a key role in accounting for the possible dependence between $\Delta_{*}$ and $Y_{*}$. We use $F_{P}$ to denote the marginal distribution function of $Y$. We use $F_{P}$ as a domain transformation function in the CIR procedure. While the CIR procedure could be performed without the use of this domain transformation, its use results in a procedure that is invariant to strictly increasing transformations of $Y$ (Westling et al., 2020). In other words, changing the scale on which response times are measured will have no impact on the estimated survival function.
We next define $\psi_{P}:=\theta_{P}\circ F_{P}^{-1}$ and corresponding primitive function $\Psi_{P}(t):=\int_{0}^{t}\psi_{P}(u)du$. The causal isotonic regression (CIR) procedure entails estimation of the parameter $\Gamma_{P}:=\Psi_{P}\circ F_{P}$, which takes the form
@@
with $Q_{P}$ the marginal distribution function of $W$ under sampling from $P$. Unlike $\theta_{0}(y_{0})$, the parameter $\Gamma_{0}(y_{0})$ is pathwise differentiable relative to the nonparametric model $\mathcal{M}$ and can therefore be estimated at $n^{-\frac{1}{2}}$–rate (Pfanzagl and Pfanzagl, 1990). Per the results of Westling et al. (2020), the efficient influence function of $P\mapsto\Gamma_{P}(y_{0})$ relative to $\mathcal{M}$ evaluated at $P_{0}$ is given by
@@
Knowledge of this efficient influence functions enables standard debiasing approaches, such as the one-step correction (Pfanzagl, 1982), to be used in estimation of $\Gamma_{0}(y_{0})$. This yields a debiased, asymptotically linear estimator $\Gamma_{n}(y_{0})$ of $\Gamma_{0}(y_{0})$, which is used in the isotonic regression procedure to construct an estimator of $\theta_{0}(y_{0})$. Our proposed procedure is as follows.
@@
    Select domain bounds $t_{0}$ and $t_{1}$ satisfying $t_{0},t_{1}]\subset(b_{0},c_{0})$.
@@
We note that, were we not to apply the domain transformation $F_{P}$, we could perform a similar procedure using the primitive parameter $y_{0}\mapsto\iint\mathbbm{1}(u\leq y_{0})\mu_{P}(u,w)Q_{P}(dw)$. In the absence of survey nonresponse, this procedure would take the same form as the isotonization method of van der Vaart and van der Laan (2006). However, it is not clear whether this approach can be framed in terms of isotonic regression using pseudo-outcomes, as described in the main text.
If for every point $y_{0}$ in the estimation domain $t_{0},t_{1}]$, either $\mu_{n}$ converges to $\mu_{0}$ or $g_{n}$ converges to $g_{0}$, Theorem 1 of Westling et al. (2020) establishes that $\theta_{n}(y_{0})$ converges in probability to $\theta_{0}(y_{0})$ for each $y_{0}\in[t_{0},t_{1}]$. The conditions used to establish distributional convergence are more complex. Roughly speaking, we require that the product of the convergence rates of $\mu_{n}$ and $g_{n}$ to their respective targets is $n^{-\frac{1}{3}}$. If these conditions are satisfied, Theorem 2 of Westling et al. (2020) establishes that, for any $y_{0}\in(b_{0},c_{0})$ with $F_{0}(y_{0})\in(0,1)$, $n^{\frac{1}{3}}\left\{\theta_{n}(y_{0})-\theta_{0}(y_{0})\right\}$ converges in distribution to $\tau_{0}(y_{0})^{\frac{1}{3}}\mathbb{W}$, where $\mathbb{W}$ follows a standard Chernoff distribution (Groeneboom and Wellner, 2001) and $\tau_{0}(y_{0}):=4\theta_{0}^{\prime}(y_{0})\kappa_{0}(y_{0})/f_{0}(y_{0})$ with
@@
A copula is a multivariate cumulative distribution function, indexed by an association parameter $\alpha$, that describes the joint distribution of two or more random variables, each with a uniform marginal distribution. Sklar’s theorem (Sklar, 1959) states that any multivariate distribution can be expressed in terms of its marginal distributions and a copula. Throughout this section, we use $F$ and $f$ to denote generic distribution and density functions, respectively, with a subscript indicating the distribution and random variable corresponding to the given distribution/density function, e.g., $F_{P;T\,|\,W}(t\,|\,w):=P(T\leq t\,|\,W=w)$.
@@
Many commonly used copula functions do not appear to yield a tractable closed form for $m_{\alpha}$ (Bernard and Czado, 2015), hindering application of the extended CIR procedure. The commonly used Frank copula (Genest, 1987) does, however, yield a usable form. The Frank copula can be written as $C_{\alpha}(u,v)=\varphi^{-1}\{\varphi(u)+\varphi(v)\}$ with generator function $\varphi:u\mapsto-\log\left\{\frac{\exp(-\alpha u)-1}{\exp(-\alpha)-1}\right\}$. The association parameter $\alpha$ takes values in $(-\infty,\infty)\setminus\{0\}$, with independence recovered as $\alpha\to 0$. The Kendall’s rank correlation coefficient $\tau$ can be expressed explicitly in terms of the Frank copula association parameter $\alpha$ via the formula $\tau=1-\frac{4}{\alpha}\left\{1-\Omega(\alpha)\right\}$, where $\Omega(\alpha)=\frac{1}{\alpha}\int_{0}^{\infty}\frac{t}{\exp(t)-1}\,dt$.
@@
$(\delta,y,w)\mapsto\$$\frac{\left(1-e^{-\alpha}\right)\mathbbm{1}(y\leq y_{0})\left\{\delta-\mu_{0}(y,w)\right\}e^{-\alpha H_{0}(y,w)}}{\alpha g_{0}(y,w)r_{\alpha}\{\mu_{0}(y,w),H_{0}(y,w)\}}-2\int_{u\leq y_{0}}\lambda_{0;\alpha}(u)F_{0}(du)$
@@
Many popular regression models used in survival analysis have been adapted for use with current status data, including the Cox proportional hazards model (Huang, 1996), the proportional odds model (Rossini and Tsiatis, 1996), and the accelerated failure time model (Rabinowitz et al., 1995). These methods are generally valid under the same assumptions on the survey response process as the adaptation of the CIR method described above. They do, however, make additional assumptions on the relationship between $T$ and $W$.
Fitting regression models using current status data is generally more complicated than with right censored data. For the Cox model, unlike in the right-censored setting, estimation of the regression coefficients cannot be completely separated from estimation of the baseline hazard function through the partial likelihood (Huang, 1996). Nonetheless, estimation and valid inference is still attainable for the regression coefficients, with the main difficulties being computational in nature. Recent work has focused on developing more efficient algorithms for computing estimates of the regression coefficients via maximum likelihood (Anderson-Bergman, 2017). Calculation of the observed Fisher information, which is used to compute analytic standard error estimates, is generally impractical. The bootstrap has been proposed as an alternative approach (Pan, 1999), although as of yet there is no theoretical proof of its validity in this context. In Section S4, we present results of a simulation study investigating the performance of the nonparametric bootstrap for construction of confidence intervals using current status data for inference on regression coefficients of a Cox model. Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance, while percentile bootstrap confidence intervals tend to be slightly anticonservative. Based on these findings, we make use of the Wald-type interval in the regression analyses we report.
As with the Cox model, the other regression models described above do not need to be altered to account for survey nonresponse, under the assumption that the regression model is correctly specified and includes all covariates that inform the response mechanism (Little and Rubin, 2019).
@@
For each simulation replicate, the aim was to estimate the distribution function of $T$ over the interval $t_{0},t_{1}]$ with $t_{0}=0.02$ (corresponding to the 5th percentile of observed response times) and $t_{1}=1.5$. The maximum follow-up time was set to $c_{0}\in\{2.1,1.8,1.65\}$. We refer to the settings corresponding to these values of $c_{0}$ as Scenarios 1, 2 and 3, respectively. Under these three choices of $c_{0}$, $t_{1}$ corresponded to the 90th, 95th, or 97.5th quantile of observed survey response times, respectively. This allowed us to explore whether performance of the complete case CIR or extended CIR procedures would degrade if $c_{0}$ was chosen such that nonresponse was more prevalent and $t_{1}$ was nearer the edge of the support of observed response times. For each value of $c_{0}$, the observed data consisted of $n$ independent replicates of $(W,Y,\Delta)$ with $Y:=\min(Y_{*},c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\mathbbm{1}(T\leq Y)$. We generated datasets of size $n\in\{500,1000,1500,2000\}$. For each simulation setting, we performed 500 simulation replicates.
The survival function was estimated (1) using only complete cases (i.e., survey respondents, or those with $Y_{*}<c_{0}$) with the existing CIR procedure described in (Westling et al., 2020), (2) using all participants with the extended CIR procedure proposed here, and (3) using the NPMLE, as implemented in the survival software package (Therneau, 2022), computed only using complete cases. For the two CIR procedures, nominal 95% confidence intervals were constructed by estimating the scale factor and using the quantiles of the Chernoff distribution. The nuisance function $\mu$ was estimated using Super Learner (van der Laan et al., 2007) with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman, 1991), generalized additive models (GAMs) (Hastie, 2017), and random forests (Wright and Ziegler, 2015). The conditional density $\pi$ was estimated using haldensify (Hejazi et al., 2022) with 10-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. For each simulation replicate, the survival function of $T$ was estimated at 100 time-points, evenly spaced on the quantile scale, between $t_{0}=0.02$ and $t_{1}=1.5$. For each setting and method, we conducted 1000 simulation replicates. Pointwise and integrated (over all 100 time-points) performance was quantified using the empirical bias and empirical confidence interval coverage.
@@
We used the extended CIR procedure, with several different combinations of nuisance estimators, to estimate the survival function of $T$. The conditional density $\pi$ was estimated using either (1) a misspecified parametric log-normal accelerated failure time model including only linear terms, or (2) haldensify, an implementation of a nonparametric approach based on the highly adaptive lasso (Hejazi et al., 2022). The outcome regression nuisance function $\mu$ was estimated using either (1) logistic regression, (2) GAMs (Hastie, 2017), (3) gradient boosted trees (Chen and Guestrin, 2016), or (4) Super Learner (van der Laan et al., 2007) with a library as described in the previous experiment. For comparison, we also included the true form for $g$ and $\mu$ (i.e., in lieu of estimating the nuisance parameters, we plugged in their known values). As long as either of the two nuisance estimators is sufficiently flexible to consistently estimate its target, we expect the extended CIR procedure to yield a consistent estimator. Furthermore, if the nuisance estimation rate is fast enough, as described in Section S1.2, we expect the resulting extended CIR estimator to not only be consistent, but to have a large-sample distribution that does not depend on the particular nuisance estimator employed. This large-sample distribution be identical as when the nuisance parameters are set to their true values. For each sample size, we conducted 500 simulations replicates; all 15 possible combinations of nuisance estimators were applied to each simulated dataset. We examined pointwise empirical bias and pointwise empirical variance. Furthermore, we examined the absolute difference between survival function estimates computed on each dataset, comparing (1) using known nuisance parameters versus using the haldensify–Super Learner combination, and (2) using the haldensify–Super Learner combination versus using the haldensify–gradient boosted trees combination. These two comparisons quantify the sensitivity of the proposed procedure to the choice of nuisance estimator when a flexible method is used in place of the ‘ideal’ method using known nuisance parameter values and when two distinct flexible methods are used, respectively.
@@
We generated datasets of size $n\in\{500,1000,2000,4000\}$. The maximum follow-up time $c_{0}$ was set to 1.65. We aimed to estimate the cumulative distribution function of $T$ over $t_{0},t_{1}]$, with $t_{0}=0.02$ and $t_{1}=1.5$. The estimation procedure was as described in Section S4.1. We compared the performance of the extended CIR procedure using the copula-based one-step estimator $\Gamma_{P_{n},\alpha}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},\alpha,y_{0}}(Z_{i})$ to the performance of the extended CIR procedure using $\Gamma_{P_{n}}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},y_{0}}(Z_{i})$, which we expect to provide consistent estimation only under conditional independence. We refer to the latter as the ‘baseline’ extended CIR. We quantified performance in terms of empirical bias. For each simulation setting, we performed 500 simulation replicates.
@@
For data analyses and simulations Super Learner (van der Laan et al., 2007) was implemented with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman, 1991), generalized additive models (Hastie, 2017), random forests (Wright and Ziegler, 2015), and gradient boosted trees (Chen and Guestrin, 2016).